# GLS: Geometry-aware 3D Language Gaussian Splatting

### v1

@@ 5.2 Indoor Surface Reconstruction
For indoor surface reconstruction, we conduct comparisons on two real-world datasets, including MuSHRoom [36] and ScanNet++ [49]. We report the metric values in Tab. 1 and Tab. 7. It can be seen that our method outperforms other 3DGS-based approaches [20, 8] without sensor depth among all metrics. When adopting sensor depth as the prior information of scene scale, our model also achieves better performance of sharpness and smoothness than DN-Splatter [40] according to the Accuracy and Normal Consistency metric. With the help of deep MLP networks, the SDF-based method MonoSDF [51] can learn the actual scale for each densely sampled 3D point. However, this method suffers from low training efficiency. Our model not only has comparable performance, but also reduces the training time by 30 times.
@@
As shown in Fig. 6, there are surface reconstruction results produced by different methods on the MuSHRoom dataset. It can be observed that DN-Splatter [40] generates severe noise on scene surfaces and even destroys object structures because of texture-less regions. On the contrary, our model reconstructs smooth and sharp scene surfaces and recovers more thin structures than ground-truth scene surfaces. This observation demonstrates the joint optimization of surface reconstruction and open-vocabulary segmentation can significantly promote the reconstruction quality. For 3DGS-based approaches without sensor depth as supervision, due to the complex camera motion and light conditions, PGSR presents unstable performance and fails in most scenes. When compared to 2DGS, our model handles shadow and high-light regions better and generates cleaner scene surfaces.
Moreover, we further evaluate the generalization ability of all methods on ScanNet++ dataset as Fig. 7 shows. These scenes encode various lighting conditions, which makes reconstructing scene surfaces more challenging. Hence, more noises occur in DN-splatter results, while our results are still sharp and smooth. MonoSDF decreases the interference of lighting conditions by over-smoothing effects, which result in the loss of several objects (e.g., the computer bracket and screen). For 3DGS-based approaches without sensor depth as supervision, PGSR and 2DGS produce meaningless results while our results successfully tackle these scenes well.
@@
Figure 7: Qualitative comparisons of indoor surface reconstruction on ScanNet++ dataset. The lighting conditions change violently in these scenes. MonoSDF loses scene details in its results.
5.3 Open-vocabulary Segmentation
@@
Open-vocabulary segmentation drops predefined labels of objects and uses texts as prompts to segment target objects. We conduct comparisons of open-vocabulary segmentation on the widely-used LERF-OVS dataset. Following Gaussian grouping [47], we employ mIoU and mBIoU as the evaluation metrics, which represent the segmentation quality of the selected object. The estimated metric scores are reported in Tab. 3. Our model outperforms state-of-the-art method OpenGaussian [43] over $4\%$ on mIoU and achieves over $20\%$ gain when compared with Gaussian grouping [47] on mBIoU. This fact illustrates that the joint optimization of two tasks improves the smoothness and sharpness of scene surfaces then makes segmentation results smoother and sharper.
Table 3: Quantitative results of open-vocabulary segmentation on LERF-OVS dataset. We follow the metrics of LangSplat [34] and OpenGaussian [43] from the paper of OpenGaussian. The best metrics are highlighted.
@@
Fig. 8 shows some open-vocabulary segmentation results of different methods. Without the optimization of surface reconstruction, OpenGaussian [43] easily treats other objects as targets and Gaussian grouping [47] exhibits noisy results. Our model successfully identifies the 3D Gaussians relevant to the query text and generates object selection with a clearer boundary.
@@ 5.4 Ablation study
We conduct sufficient ablation experiments to study the effect of different regularization terms, by disabling each one and enabling others. The results of indoor surface reconstruction and open-vocabulary segmentation are reported in Tab. 4 and Tab. 5 respectively.
Table 4: Ablation study of indoor surface reconstruction on a scene of MuSHRoom dataset. The best metrics are highlighted.
@@
No $\mathcal{L}_{n}$ 0.0864 0.0758 0.0811 0.7390 0.5559
No $\mathcal{L}_{n}$, No $\mathcal{L}_{d}$ 0.1031 0.0828 0.0929 0.7208 0.5247
No $\mathcal{L}_{s}$ 0.0956 0.0852 0.0904 0.8035 0.5453
No $\mathcal{L}_{clip}$ 0.1037 0.0985 0.1011 0.8113 0.5434
No $\mathcal{L}_{m}$ 0.0731 0.0871 0.0801 0.8343 0.5695
All 0.0720 0.0828 0.0774 0.8267 0.5784
Table 5: Ablation study of open-vocabulary segmentation in on a scene of LERF-OVS dataset. The best metrics are highlighted.
@@
No $\mathcal{L}_{n}$ 30.80 18.99
No $\mathcal{L}_{d}$ 28.12 19.76
No $\mathcal{L}_{s}$ 30.84 20.87
No $\mathcal{L}_{clip}$ 23.20 15.27
No $\mathcal{L}_{m}$ 26.78 16.15
All 32.15 23.93
Figure 8: Qualitative comparisons of open-vocabulary segmentation on LERF-OVS dataset, between state-of-the-art methods (OpenGaussian [43] and Gaussian grouping [47]) and our model.
@@
The Normal prior supplies the regions with smooth surfaces, where shadow and highlight exist in an indoor scene. For indoor surface reconstruction, $\mathcal{L}_{n}$ remarkably improves surface smoothness in terms of the Normal Consistency metric. For open-vocabulary segmentation, $\mathcal{L}_{n}$ helps the full model to produce accurate boundary of segmentation results by the joint optimization according to the metric mBIoU metric.
@@
Fig. 5 illustrates the visual effects of $\mathcal{L}_{d}$ for indoor surface reconstruction. For open-vocabulary segmentation, the performance of our model degrades when disabling $\mathcal{L}_{d}$. Although $\mathcal{L}_{d}$ is designed for refining the unbiased depth, it also strengthens the quality of segmentation results through joint optimization.
@@
To determine whether $\mathcal{L}_{s}$ is necessary for two tasks, we disable it from our full model. $\mathcal{L}_{s}$ enhances the performance of our model among all metrics, by connecting the surface normal estimated from the unbiased depth and Gaussian semantic features.
@@
Then we disable $\mathcal{L}_{clip}$, to verify it is a significant regularization term for two tasks. We introduce CLIP features to strengthen the representation ability of Gaussian semantic features, and which plays the most important role in open-vocabulary segmentation. It also significantly maintains the sharpness of reconstructed indoor surfaces via joint optimization.
@@
We further explore the effect of $\mathcal{L}_{m}$ by dropping the segmentation results of DEVA as supervision. As Tab. 5 reports, $\mathcal{L}_{m}$ faithfully increases the completeness and sharpness of segmentation results.
5.5 Additional Results
Surface reconstruction on LERF-OVS dataset.
Novel view synthesis.
Finally, we evaluate the performance of our model in the task of novel view synthesis. The results are presented in Tab. 6. Our model achieves comparable performance against the state-of-the-art methods.
Table 6: Quantitative results of novel view synthesis on ScanNet++ dataset. The best metrics are highlighted.
@@ 6 Conclusion
In this work, we present GLS, a novel 3DGS-based framework that effectively combines indoor surface reconstruction and open-vocabulary segmentation. We propose leveraging 2D geometric and semantic cues to optimize the performance of 3DGS on two tasks jointly. We design two novel regularization terms to enhance the sharpness and smoothness of the scene surface, and then improve the segmentation quality. Comprehensive experiments on both open-vocabulary segmentation and indoor surface reconstruction tasks illustrate that GLS outperforms state-of-the-art methods quantitatively and qualitatively. Besides, the ablation study explores the effectiveness of each regularization term on two tasks.
@@
The proposed method follows the natural limitation of TSDF fusion, whose completeness relies on the number of captured views. Our model generates empty geometry of selected objects in unseen views of the scene. We will introduce image-to-3D models to supply unseen information in future work.
Supplementary Material
7 Additional Algorithm Details
@@
8 Additional Experimental Details

### v2

@@ 5.2 Indoor Surface Reconstruction
For indoor surface reconstruction, we conduct comparisons on two real-world datasets, including MuSHRoom[41] and ScanNet++[54]. We report the metric values in Tab. 1 and Tab. 7. It can be seen that our method outperforms other 3DGS-based approaches[24, 10] without sensor depth among all metrics. When adopting sensor depth as the prior information of scene scale, our model also achieves better performance of sharpness and smoothness than DN-Splatter[46] according to the Accuracy and Normal Consistency metric. Compared with the state-of-the-art method FDS[12], our results have better smoothness, accuracy and training efficiency.
@@
As shown in Fig. 6, there are surface reconstruction results produced by different methods on the MuSHRoom dataset. It can be observed that DN-Splatter[46] generates severe noise on scene surfaces and even destroys object structures because of texture-less regions. On the contrary, our model reconstructs smooth and sharp scene surfaces and recovers more thin structures than ground-truth scene surfaces. This observation demonstrates the joint optimization of surface reconstruction and 3D open-vocabulary segmentation can significantly promote the reconstruction quality. For 3DGS-based approaches without sensor depth as supervision, due to the complex camera motion and light conditions, PGSR presents unstable performance and fails in most scenes. When compared to 2DGS, our model handles shadow and high-light regions better and generates cleaner scene surfaces.
Moreover, we further evaluate the generalization ability of all methods on ScanNet++ dataset as Fig. 7 shows. These scenes encode various lighting conditions, which makes reconstructing scene surfaces more challenging. Hence, more noises occur in DN-splatter results, while our results are still sharp and smooth. For 3DGS-based approaches without sensor depth as supervision, PGSR and 2DGS produce meaningless results while our results successfully tackle these scenes well.
@@
Figure 7: Qualitative comparisons of indoor surface reconstruction on ScanNet++ dataset. The lighting conditions change violently in these scenes.
5.3 3D open-vocabulary segmentation
@@
3D open-vocabulary segmentation drops predefined labels of objects and uses texts as prompts to segment target objects. We conduct comparisons of 3D open-vocabulary segmentation on the widely-used LERF-OVS dataset. Following Gaussian grouping[52], we employ mIoU and mBIoU as the evaluation metrics, which represent the segmentation quality of the selected object. The estimated metric scores are reported in Tab. 3. Our model outperforms state-of-the-art method OpenGaussian[49] over $4\%$ on mIoU and achieves over $20\%$ gain when compared with Gaussian grouping[52] on mBIoU. This fact illustrates that the joint optimization of two tasks improves the smoothness and sharpness of scene surfaces then makes segmentation results smoother and sharper.
Table 3: Quantitative results of 3D open-vocabulary segmentation on LERF-OVS dataset. We follow the metrics of LangSplat[39] and OpenGaussian[49] from the paper of OpenGaussian. The best metrics are highlighted.
@@
Fig. 8 shows some 3D open-vocabulary segmentation results of different methods. Without the optimization of surface reconstruction, OpenGaussian[49] easily treats other objects as targets and Gaussian grouping[52] exhibits noisy results. Our model successfully identifies the 3D Gaussians relevant to the query text and generates object selection with a clearer boundary.
@@ 5.4 Ablation study
We conduct sufficient ablation experiments to study the effect of different regularization terms, by disabling each one and enabling others. The results of indoor surface reconstruction and 3D open-vocabulary segmentation are reported in Tab. 4 and Tab. 5 respectively.
Table 4: Ablation study of indoor surface reconstruction without the sensor depth. The best metrics are highlighted. The red box means that the loss function is designed for 3D open-vocabulary segmentation.
@@
No $\mathcal{L}_{n}$ 0.1850 0.6893 0.2322
No $\mathcal{L}_{d}$ 0.1077 0.8023 0.4440
No $\mathcal{L}_{s}$ 0.1303 0.7910 0.4356
No $\mathcal{L}_{clip}$ 0.0966 0.8393 0.4614
No $\mathcal{L}_{m}$ 0.0951 0.8425 0.4671
All 0.0814 0.8474 0.5127
Table 5: Ablation study of 3D open-vocabulary segmentation. The best metrics are highlighted. The red box means that the loss function is designed for indoor surface reconstruction.
@@
No $\mathcal{L}_{n}$ 32.90 27.61
No $\mathcal{L}_{d}$ 31.81 27.37
No $\mathcal{L}_{s}$ 33.98 29.08
No $\mathcal{L}_{clip}$ 29.14 25.01
No $\mathcal{L}_{m}$ 28.71 22.62
All 42.78 38.19
Figure 8: Qualitative comparisons of 3D open-vocabulary segmentation on LERF-OVS dataset, between state-of-the-art methods (OpenGaussian[49] and Gaussian grouping[52]) and our model.
@@
The Normal prior supplies the regions with smooth surfaces, where shadow and highlight exist in an indoor scene. For indoor surface reconstruction, $\mathcal{L}_{n}$ remarkably improves surface smoothness in terms of the Normal Consistency metric. For 3D open-vocabulary segmentation, $\mathcal{L}_{n}$ helps the full model to produce accurate boundary of segmentation results by the joint optimization according to the metric mBIoU.
@@
Fig. 5 illustrates the visual effects of $\mathcal{L}_{d}$ for indoor surface reconstruction. For 3D open-vocabulary segmentation, the performance of our model degrades when disabling $\mathcal{L}_{d}$. Although $\mathcal{L}_{d}$ is designed for refining the unbiased depth, it also strengthens the quality of segmentation results through joint optimization.
@@
To determine whether $\mathcal{L}_{s}$ is necessary for two tasks, we disable it from our full model. $\mathcal{L}_{s}$ enhances the performance of our model among all metrics, by connecting the surface normal estimated from the unbiased depth and CLIP features.
@@
Then we disable $\mathcal{L}_{clip}$, to verify it is a significant regularization term for two tasks. We introduce CLIP features to strengthen the representation ability of Gaussian semantic features, and which plays the most important role in 3D open-vocabulary segmentation. It also significantly maintains the smoothness and sharpness of reconstructed indoor surfaces via joint optimization.
@@
We further explore the effect of $\mathcal{L}_{m}$ by dropping the segmentation results of DEVA as supervision. As Tab. 5 reports, $\mathcal{L}_{m}$ faithfully increases the sharpness of rexonstruction and segmentation results.
5.5 Novel view synthesis
Finally, we evaluate the performance of our model in the task of novel view synthesis. The results are presented in Tab. 6. Our model achieves comparable performance with DN-Splatter, and better performance against other geometry-aware approaches in the field of indoor surface reconstruction. We also compare our method with OpenGaussian in the field of 3D open-vocabulary segmentation, our method renders more accurate novel views than it.
Table 6: Quantitative results of novel view synthesis on ScanNet++ and LERF-OVS dataset. The best metrics are highlighted.
@@ 6 Conclusion
In this work, we present GLS, a novel 3DGS-based framework that effectively combines indoor surface reconstruction and 3D open-vocabulary segmentation. We propose leveraging 2D geometric and semantic cues to optimize the performance of 3DGS on two tasks jointly. We design two novel regularization terms to enhance the sharpness and smoothness of the scene surface, and then improve the segmentation quality. Comprehensive experiments on both 3D open-vocabulary segmentation and indoor surface reconstruction tasks illustrate that GLS outperforms state-of-the-art methods quantitatively and qualitatively. Besides, the ablation study explores the effectiveness of each regularization term on two tasks.
@@
The proposed method follows the natural limitation of TSDF fusion, whose completeness relies on the number of captured views. Our model generates empty geometry of selected objects in unseen views of the scene. Introducing image-to-3D models to supply unseen information can help solve this issue.
Appendix A Q&A
  • Q1.
    Please explain the benefits of the proposed joint optimization.
  • A2.
    We have analyzed the benefits of the proposed joint optimization of indoor surface reconstruction and 3D open-vocabulary segmentation from Line_037 to Line_056 of the manuscript. Theoretically, the optimization goals of the two tasks are the smoothness and sharpness of generated results (including the surface mesh and the segmentation mask). Then we can optimize two tasks simultaneously through the gradient descent strategy. The marked metrics in Table 4 and Table 5 of the manuscript demonstrate the effectiveness of the proposed joint optimization in two fields. The qualitative results of the ablation study are shown in  Fig. 11. The performance of GLS degenerates reasonably when we disable a regularization term and enable others.
  • Q3.
    Why does GLS only focus on indoor scenes?
  • A4.
    Considering the unique challenges of indoor scenes and the limitation of geometric and semantic cues in outdoor scenes, there are two main reasons as follows:
    • i.
      For surface reconstruction, we follow DN-Splatter [46] and FDS [12], then focus on achieving high-fidelity reconstruction in indoor scenes. Unlike outdoor scenes (e.g., the Tanks and Temples dataset [32]), indoor scenes suffer from more challenging conditions. On the one hand, indoor scenes are always captured by free camera trajectories, which causes the imbalanced allocation of spatial representation capacity. On the other hand, there are a lot of texture-less regions (such as white walls and floors) in indoor scenes. These issues cause extreme ambiguity of multi-view reconstruction and large performance degeneration of the state-of-the-art method PGSR [10]. However, GLS tackles these issues well because of the robust joint optimization.
    • ii.
      As Fig. 14 shows, geometric cues predicted from the pre-trained monocular normal model, suffer from inconsistent noises across different views in outdoor scenes. The semantic cues only consist of a few instances and encode large ambiguity of object structures. These issues are harmful to the joint optimization of GLS.
      Hence, we focus on the surface indoor scenes where the quality of geometric cues is more stable and semantic distribution is highly predictable. To illustrate the generalization ability of GLS, We also evaluate the performance of our model in a large-scale indoor scene (‘Meetingroom’) of the Tanks and Temples dataset, the result is presented in  Fig. 10. Our method successfully handles the high-light and dark regions and reconstructs a smoother scene surface than PGSR [10].
  • Q5.
    How about the robustness of GLS under weak semantic priors?
  • A6.
    Most view-consistent segmentation priors we used on the MuSHRoom dataset are noisy or incomplete, such as  Fig. 13 shows. It demonstrates that our method is certainly robust. We also evaluate the reconstruction result under different quality of CLIP features, by disabling the denoising procedure of pre-processing to obtain the low-quality CLIP features. The result is shown in  Fig. 12. High-quality CLIP features reduce shadow interference. This result also illustrates the effectiveness of the proposed joint optimization.
  • Q7.
    How about applications of GLS?
  • A8.
    Thanks to the high-quality surface reconstruction and segmentation results of GLS, we develop several tools to explore applications of GLS:
    • i.
      GLS + Blender: With the help of Blender [16], we can achieve the scene-editing effects, including moving the target object and adding objects.
    • ii.
      GLS + nerfview: We develop an interactive tool based on nerfview [1], to show the scene geometry (depth and normal) and open-vocabulary attention in free views.
    • iii.
      GLS + nerfview + GPT-4V: As  Fig. 9 shows, we also add the GPT-4V [2] to ‘ii’ to achieve intelligent interaction. Given a user’s request, our tool can find the object that solves the request and extract the scene geometry used to help the user plan the path to the target object. We believe this tool has potential value in the field of embodied intelligence.
    Our supplementary videos consist of two application demos (‘application_i.mp4’ and ‘application_ii.mp4’). We will release all tools when this paper is accepted.
Appendix B Additional Algorithm Details
@@
Figure 12: Effect of different quality of CLIP features.
Figure 13: Example of the view-consistent segmentation results of DEVA [11]. We directly use these results as the supervision of the rendered mask.
Figure 14: Limitation of geometric and semantic cues in outdoor scenes. The geometric cues predicted from the pretrained monocular normal model, suffer from inconsistent noises across different views. The semantic cues encode large ambiguity of object structures.
@@
Appendix C Additional Experimental Details
@@ Appendix D Additional results
Appendix E Future Work
In the future, we have two plans:
  • i.
    To address the limitation of TSDF Fusion, we can first segment object mesh based on GLS to supply the object scale and orientation, then leverage the image-to-3D models to obtain the whole object body from the segmented object appearance.
  • ii.
    To extend GLS to outdoor scenes, we can follow VCR-GauS [11] to estimate the confidence of geometric and semantic cues to solve their issues.